In an exciting game, a baseball player manages to safely slide into second base. The mass of the baseball player is 85.7 kg and the coefficient of kinetic friction between the ground and the player is 0.44.

(a) Find the magnitude of the frictional force in newtons. 369.96 N Correct

(b) It takes the player 1.8 s to come to rest. What was his initial velocity (in m/s)? m/s
Answer #1

Part A.

Frictional Force is given by:

Ff = -uk*N = -uk*m*g

Ff = -0.44*85.7*9.81 = -369.92 N

Magnitude of friction force = |Ff| = 369.92 N

Part B.

Using 1st kinematic equation:

V = U + a*t

V = final velocity = 0 m/sec

t = time taken = 1.8 sec

a = acceleration = -uk*g = -0.44*9.81 m/sec^2

(from Newton's 2nd law: Fnet = m*a = -uk*m*g, which gives, a = -uk*g)

So,

U = V - a*t

U = 0 - (-0.44*9.81)*1.8

U = 7.77 m/sec = Initial Velocity of player

Please Upvote.